OpenStudy (anonymous):

f(x)=sqrt((ln6x+7)(7x−3))

OpenStudy (anonymous):

thank you for letting sam to have madel

OpenStudy (anonymous):

f(x)=(X^2+9)^5/(x^2+6)^2

OpenStudy (unklerhaukus):

what is the question here?

OpenStudy (anonymous):

logarithmic differentiation

OpenStudy (unklerhaukus):

which function?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

can you post a screenshot please?

jimthompson5910 (jim_thompson5910):

the problem is a bit ambiguous

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Ok sure

jimthompson5910 (jim_thompson5910):

wait, its nowhere close to sqrt((ln6x+7)(7x−3)) are you sure this is the right one?

OpenStudy (anonymous):

yes this question and root question

jimthompson5910 (jim_thompson5910):

well the second one was clearer than the first, so no need for a screenshot (in my opinion) for the second one

OpenStudy (anonymous):

ok, sorry about that

jimthompson5910 (jim_thompson5910):

ok so it's completely flipped around, thought so, but wasn't sure

jimthompson5910 (jim_thompson5910):

you would use the chain rule twice, then the product rule once you hit the product

OpenStudy (anonymous):

for the root one?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

I see! what part do i use chain rule?

jimthompson5910 (jim_thompson5910):

for the ln( ... ) part and the sqrt( ... ) part

OpenStudy (anonymous):

(6x+7)^1/2 (7x-3)^1/2?

jimthompson5910 (jim_thompson5910):

first you would take the derivative of ln(x) to get 1/x but x is some expression with a square root of a product

jimthompson5910 (jim_thompson5910):

so it's really 1/(that expression) but then you use the chain rule to multiply it by the derivative of that expression

jimthompson5910 (jim_thompson5910):

then you would use the chain rule again to derive (6x+7)(7x-3)

jimthompson5910 (jim_thompson5910):

which you would use the product rule at this point

jimthompson5910 (jim_thompson5910):

so as you can see, it gets ugly fast

OpenStudy (anonymous):

in(6x+7)(in7x-3)?

jimthompson5910 (jim_thompson5910):

it's L n (lower case L though)

jimthompson5910 (jim_thompson5910):

and that seems too simple to be the final answer

OpenStudy (anonymous):

no the first step, haha

jimthompson5910 (jim_thompson5910):

oh, no that's not correct either

OpenStudy (anonymous):

oh :(

jimthompson5910 (jim_thompson5910):

the first step is to derive ln( sqrt( (6x+7)(7x-3) ) ) to get 1/( sqrt( (6x+7)(7x-3) ) ) * d/dx[ sqrt( (6x+7)(7x-3) ) ]

jimthompson5910 (jim_thompson5910):

you would then take the derivative of sqrt( (6x+7)(7x-3) )

OpenStudy (anonymous):

take?

OpenStudy (anonymous):

(6+7)'7x-3)^1/2?

jimthompson5910 (jim_thompson5910):

the derivative of sqrt( (6x+7)(7x-3) ) is _____

jimthompson5910 (jim_thompson5910):

what you got isn't quite it

OpenStudy (anonymous):

In(6)(7)?

jimthompson5910 (jim_thompson5910):

no just focus on sqrt( (6x+7)(7x-3) ) for now

OpenStudy (anonymous):

ok, sqrt is always confusing me

jimthompson5910 (jim_thompson5910):

if y = sqrt(x) then y ' = 1/(2*sqrt(x))

OpenStudy (anonymous):

Oh! 1/ (6x+7)(7x-3) )

OpenStudy (raden):

well, before i derive it. try to expanding (6x+7)(7x-3)=42x^2+31x-21 so, f(x) = ln{sqrt(42x^2+31x-21)} now, we derive it by chain rule a re-explain what is @jim_thompson5910 said before, if given y = sqrt(f(x)) then y ' = f '/2sqrt(f(x)) and we knowed that the derivative of ln(u) = 1/u thus, the derivative of y can be y ' = {(42x^2+31x-21)'/2sqrt(42x^2+31x-21)}/(sqrt42x^2+31x-21) y ' = (84x+31)/2(42x^2+31x-21)

OpenStudy (raden):

look not enough good to read :) i will figure out that :